# Fine-Tuning Flow Matching via Maximum Likelihood Estimation of Reconstructions

### v1

@@ Appendix D Experiment Details and Additional Results
The experiments supported several execution modes: FM pretraining mode; MLE fine-tuning mode or residual fine-tuning with a specialized network architecture; and testing mode with 10 test episodes, and 50 runs per episode by default. The maximum number of steps was automatically configured according to the environment settings. This modular configuration supports high customizability while ensuring the reproducibility of all experimental results.
Table D.1 shows the hyperparameters used in FM and diffusion policy. Table D.2 shows the task summary. Our evaluations were all conducted within the MuJoCo or Gym simulation environment. Specifically, the Push-T simulation environment is developed using Gym, whereas the Franka Kitchen and Robomimic environments utilize MuJoCo for simulation.
@@
Table D.1: Hyperparameters for FM and diffusion policy. Ta: action horizon. Tp: action prediction horizon. ObsRes: environment observation resolution. D-Net: diffusion/flow matching network. D-Par: diffusion/flow matching network number of parameters in millions. V-Enc: vision encoder. V-Par: vision encoder number of parameters in millions. P-Lr: learning rate in pretraining. F-Lr: learning rate in MLE fine-tuning.
@@
Table D.2: Tasks Summary. Rob: number of robots. Obj: number of objects. ActD: action dimension. PH: proficient-human demonstration. Steps: max number of rollout steps. Franka Kitchen and Robomimic involve 6D robot and gripper actions in the joint space. Push-T focuses on robot end-effector trajectories.
Appendix E The Use of Large Language Models
We employed some Large Language Models (LLMs) to polish the writing and identify grammatical errors.
@@ Appendix E The Use of Large Language Models
Figure E.1: Different Simulation Environments in MuJoCo or Gym.
@@
(b) Images generated by FM at $3.5\times 10^{5}$ step, FID: 3.61.
@@
Figure E.2: Visualization Results of generated images under the same initial noise points.

### v2

@@ C.2.1 Experimental Details
Table 3 shows the hyperparameters used in FM and diffusion policy. Table 4 shows the task summary. Our evaluations were all conducted within the MuJoCo or Gym simulation environment. Specifically, the Push-T simulation environment is developed using Gym, whereas the Franka Kitchen and Robomimic environments utilize MuJoCo for simulation.
@@
Table 3: Hyperparameters for FM and diffusion policy. Ta: action horizon. Tp: action prediction horizon. ObsRes: environment observation resolution. D-Net: diffusion/flow matching network. D-Par: diffusion/flow matching network number of parameters in millions. V-Enc: vision encoder. V-Par: vision encoder number of parameters in millions. P-Lr: learning rate in pretraining. F-Lr: learning rate in MLE fine-tuning.
@@
Table 4: Tasks Summary. Rob: number of robots. Obj: number of objects. ActD: action dimension. PH: proficient-human demonstration. Steps: max number of rollout steps. Franka Kitchen and Robomimic involve 6D robot and gripper actions in the joint space. Push-T focuses on robot end-effector trajectories.
@@
Figure C.3: Different Simulation Environments in MuJoCo and Gym.
C.2.2 Additional Results
Figure C.4 presents the loss curves corresponding to the results shown in Figure 5(a). It can be observed that, compared to ERA5 (Figure 5(b)), the MLE fine-tuning training error curve for the robot policy is smoother, indicating a relatively lower training difficulty. This is precisely because the generated variable dimension of the robot policy is relatively low (11-dimensional in the Franka Kitchen environment), which reduces the difficulty of ODE simulation. However, compared to FM pre-training, the training difficulty is still somewhat higher. In the fourth subplot (where fine-tuning starts at 50 epochs), after going through a phase of optimization plateau, the training error declines rapidly toward the end, corresponding to the relatively noticeable rise in the final segment of the success rate curve shown in Figure 5(a).
Figure C.4: Training Loss curves of MLE Fine-tuned Models at Different Stages in the Franka Kitchen Environment.
C.3 Meteorological Data
C.3.1 Experimental Details
The experimental evaluation employs the Flow Neural Process (FNP) model on ERA5 meteorological reanalysis data.
ERA5 data for the US region are normalized and preprocessed with circular handling, providing 950–1000 randomly sampled context points per instance. The FNP model architecture incorporates a continuous normalizing flow with Transformer-based attention mechanisms and high-frequency positional encoding.
Base training runs for 200 epochs with a batch size of 100, learning rate $5\times 10^{-4}$, and cosine annealing. Validation occurs every 20 epochs, and checkpoints are saved accordingly. The MLE fine-tuning phase loads the best base checkpoint and continues for 100 epochs with a reduced learning rate $1\times 10^{-5}$, batch size 32, weight decay $1\times 10^{-7}$, and 16 flow-matching sampling steps. Fine-tuning employs distributed training across five GPUs with linear warmup and constant scheduling. During residual fine-tuning, we retain the pre-training normalization architecture but with a reduced scale (fewer layers and parameters) and zero-initialized state input to ensure decoupling. In contrast, the ablation study replaces the GAV module with a simpler, single-hidden-layer MLP (128 nodes). For complete implementation specifics, consult the source code. Evaluation uses 50 Monte Carlo samples to compute log-likelihood, mean squared error, and uncertainty calibration.
Baseline Models include Conditional Neural Process (CNP) (Garnelo et al., 2018a), Neural Process (NP) (Garnelo et al., 2018b), attention mechanisms Neural Process (ANP) (Kim et al., 2019) and Transformer Neural Process (TNP) (Nguyen and Grover, 2022). The implementation employs Nguyen and Grover (2022); Lee et al. (2020).
C.3.2 Additional Results
Figure C.5 presents the visualization results of an experiment. The first four subplots depict temperature, while the last one illustrates wind speed. Figure C.6 presents the loss curves corresponding to the results shown in Figure 5(b).
Figure C.5: Visualization of predicted results in ERA5.
Figure C.6: Training Loss curves of MLE Fine-tuned Models at Different Stages in the ERA5 dataset.
C.4 Image Generation
We perform an experiment on unconditional CIFAR-10 generation from a Gaussian source to examine how OT-CFM performs in the high-dimensional image setting. We use a similar setup to that of (Lipman et al., 2023) and (Tong et al., 2024), including the time-dependent U-Net architecture from Nichol and Dhariwal (2021) that is commonly used in diffusion models. We choose channels =128=128, depth =2=2, channels multiple =[1,2,2,2]=[1,2,2,2], heads =4=4, heads channels =64=64, attention resolution =16=16, dropout =0.1=0.1, batch size per gpu =128=128, gpus =4=4, epochs =2000=2000. We use a constant learning rate, set to $2\times 10^{-4}$ in the pre-training stage and $5\times 10^{-6}$ in the fine-tuning stage. To prevent training instabilities and variance, we clip the gradient norm to 1 and rely on exponential moving average with a decay of 0.99.
The baseline models are chosen as DDPM (Ho et al., 2020), FM with Variance Exploding (VE) path and Optimal Transport (OT) path (Lipman et al., 2023), and Stochastic Interpolants (SI) (Albergo et al., 2023). We train our fine-tuned FM and report the Fréchet inception distance (FID) in Table 5.
From Table 5, we can see that fine-tuning FM can improve the performance on FID with almost the same number of function evaluations (NFE). These demonstrate the significance of fine-tuning FM.
Table 5: FID score and number of function evaluations (NFE) for different ODE solvers: fixed-step Euler integration with 100 and 1000 steps and adaptive integration (Hairer et al., 1993, dopri5). The adaptive solver is significantly better than the Euler solver in fewer steps. First results are from (Lipman et al., 2023) and the next three from (Tong et al., 2024). The last rows report the results of our fine-tuned FM. NFE / sample $\rightarrow$ 100 1000 Adaptive Algorithm $\downarrow$ FID FID FID NFE DDPM 7.48 274 VP-FM 7.772 4.048 4.335 525.92 OT-FM 4.640 3.822 3.655 143.00 S.I. 4.488 4.132 4.009 146.12 I-CFM 4.461 3.643 3.659 146.42 OT-CFM 4.443 3.741 3.577 133.94 FT-FM (ours) 4.451 3.620 3.496 146.42
The visualization results of the generated images are presented in Figure C.7. We can see that fine-tuning induces minimal alteration to the latent space semantics. Images generated from identical noise points retain consistent content, with a slight improvement in reconstruction fidelity. From Table 5, we can see that fine-tuning FM can slightly improve the performance on FID with almost the same NFE. The effectiveness on image data, however, is less pronounced compared to the results on policy and other tasks presented previously. This is because the generative variables for images are high-dimensional ($32\times 32\times 3=3071$ in this case), and pixels exhibit strong, complex correlations. This significantly increases the difficulty of the required MLE training. In future work, we plan to conduct generation experiments within the compressed latent space.
@@ C.4 Image Generation
(b) Images generated by FM at $3.5\times 10^{5}$ step, FID: 3.57.
@@
Figure C.7: Visualization Results of generated images under the same initial noise points.